OpenStudy (taylor.joe):

Help Please!!! if f(x)= 3x+3 ,here f^-1 (x) = ____

OpenStudy (foxy7029):

need help

OpenStudy (taylor.joe):

I need help yes please.

OpenStudy (foxy7029):

ok

OpenStudy (foxy7029):

try -3-3

OpenStudy (foxy7029):

i think

OpenStudy (taylor.joe):

So just answer -3-3?

OpenStudy (taylor.joe):

I need the inverse of the function i think

OpenStudy (foxy7029):

ok ummmm

OpenStudy (foxy7029):

Parth help him i'm thinking

Parth (parthkohli):

Aight. Whenever you have to find the inverse of a function, you replace \(x\) with \(y\).\[y = 3x+3\]\[x \to y, y \to x\]\[x = 3y + 3\]now isolate \(y\).\[x-3 = 3y\]\[y = \frac{x-3}3\]what's the inverse then?

OpenStudy (taylor.joe):

is it 3+3y?

OpenStudy (taylor.joe):

Im sorry i really dont get this stuff.